OpenStudy (anonymous):

Find the limit as x approaches negative infinity of (x+sqrt(x^2 + x +1)) ? Sorry to sound needy, but HEEEEEEEEELP.

OpenStudy (anonymous):

\[\lim_{x\to-\infty}\left(x+\sqrt{x^2+x+1}\right)\] \[\begin{align*}x+\sqrt{x^2+x+1}\cdot\frac{x-\sqrt{x^2+x+1}}{x-\sqrt{x^2+x+1}}&=\frac{x^2-\left(x^2+x+1\right)}{x-\sqrt{x^2+x+1}}\\ &=\frac{-x-1}{x-\sqrt{x^2+x+1}}\\ &=\frac{-x-1}{x-\sqrt{x^2}\sqrt{1+\frac{1}{x}+\frac{1}{x^2} }}\\ &=\frac{-x-1}{x-|x|\sqrt{1+\frac{1}{x}+\frac{1}{x^2} }} \end{align*}\] Since \(x\to-\infty\), you have \(|x|=-x\): \[\begin{align*}\frac{-x-1}{x-(-x)\sqrt{1+\frac{1}{x}+\frac{1}{x^2} }}&=\frac{-x-1}{x+x\sqrt{1+\frac{1}{x}+\frac{1}{x^2} }}\\ &=\frac{-1-\frac{1}{x}}{1+\sqrt{1+\frac{1}{x}+\frac{1}{x^2} }} \end{align*}\] Taking the limit should be easy now.

OpenStudy (anonymous):

You again! Thanks. :D

OpenStudy (anonymous):

You're welcome : )